OpenStudy (anonymous):

Rationalize √2+√5 -------- √2-√5

OpenStudy (anonymous):

You need to multiply the top and bottom by sqrt(2) + sqrt(5) (sqrt(2) + sqrt(5))(sqrt(2) + sqrt(5)) = 2 + sqrt(10) + sqrt(10) + 5 = 7 + 2sqrt(10) (sqrt(2) - sqrt(5))(sqrt(2) + sqrt(5)) = 4 - 5 = -1 7 + 2sqrt(10) ----------- -1 -7 - 2sqrt(10) is the answer

OpenStudy (anonymous):

Thank You